a 20 foot ladder leaning against the building reches a point 16 feet above the ground. how far is the foot of the ladder from the foot of the building?
Answer by rfer(16322)

a^2+b^2=c^2
16^2+b^2=20^2
256+b^2=400
b^2=400-256
b^2=144
sqrt b^2=12ft